• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Keller

    Log entry time 03:31:18 on March 15, 2012

    Entry number 366540

    keyword=[Target] supply valve

    LHe level in magnet at 100% again. The EV91102 value between buffer dewar and magnet seems to be stuck at 2.99 (according to the batch fill readout). We are still waiting for tuned beam, may try to heat up valve when an if buffer drops and/or MCC is finished. Buffer dewar presently at 70%.